OpenStudy (abbycross167):

Graph the following equations using the point-slope method. Graph the point from the equation in yellow, your other points in red and your line in blue. y + 3 = (1/4)(x - 2)

OpenStudy (abbycross167):

@RAM231

OpenStudy (ram231):

you basically have to do the same thing as we did in the other problem. distribute the (1/4)(x-2) so it would be 1/4x - (1/4)x-2 what would that second set of numbers equal. when you distribute this numbers

OpenStudy (abbycross167):

umm I don't know

OpenStudy (ram231):

wait. I mean (1/4) multiplied by -2

OpenStudy (abbycross167):

so -1/2?

OpenStudy (ram231):

if you change the 1/4 into a decimal, it would be 0.25. like quarters right? Wouldn't it be easier to change the fraction into a decimal to figure out the answer faster?

OpenStudy (abbycross167):

yes ma'am

OpenStudy (ram231):

so what would be the answer to 0.25 multiplied by -2?

OpenStudy (abbycross167):

so -0.5?

OpenStudy (ram231):

yeah. so now the equation would be y+3=-0.5 how do you get the 3 to the other side?

OpenStudy (abbycross167):

-0.5+3=2.5

OpenStudy (ram231):

I would actually do y+3=-0.5... I would subtract the into get y=-0.5-3 -3 would be the y intercept and -0.6 would be m in the equation mx+b=c2 I mean the x isn't there. but it still is kinda there.

OpenStudy (abbycross167):

ok great thank you soooo much again!!!

OpenStudy (ram231):

yeah. what program do you do for school?

OpenStudy (abbycross167):

I'm in algebra 1A